-4(7-2x) = 3(2x-1) - 5

Answer by checkley71(8403)

-4(7-2X)=3(2X-1)-5
-28+8X=6X-3-5
8X-6X=-8+28
2X=20
X=20/2
X=10 ANSWER.
PROOF
-4(7-2*10)=3(2*10-1)-5
-4(7-20)=3(20-1)-5
-4*-13=3*19-5
52=57-5
52=52